what is the missing number? 7x6=(7x3)+(7x_)
Answer #1

Note that 7*6 - 7*3 =

= 7*(6-3) = 7*3

Hence the missing number is 3.

Also we can do like this: 7*6 = 42, 7*3 = 21

42-21 = 21 = 7*3